# HumanEval-XL: A Multilingual Code Generation Benchmark for Cross-lingual Natural Language Generalization

### 26 Feb 2024 v1

@@ 1.   Introduction
Large language models (LLMs) have made significant progress in code generation, with a predominant focus on the evaluation of generating Python code within an English-centric context (Chen et al., 2021; Austin et al., 2021; Hendrycks et al., 2021; Nijkamp et al., 2023b). In recent years, there has been a growing interest in expanding beyond English-Python code generation to English-multilingual programming languages (PLs) (Athiwaratkun et al., 2023). Some works have also made efforts to incorporate different natural languages (NLs), such as Spanish and Russian, although their coverage of different natural languages remains rather limited (Wang et al., 2022; Chai et al., 2022; Scao et al., 2022). A noticeable gap in these benchmarks is their limited ability to provide a rigorous assessment of multilingual NL to code generation. This overlooks the wide range of different NLs, raising important questions to the evaluation of cross-lingual NL generalization of LLMs. It further underscores the need for a more comprehensive evaluation framework that addresses the gap and captures the diverse linguistic landscape in code generation.
@@ 2.   Related work
LLMs undergo a critical phase of pre-training on a rich collection of resources that include both multilingual PL data and NL data. Some studies have delved into the intricacies of LLMs in multiple PLs (Fried et al., 2023; Wang et al., 2023; Nijkamp et al., 2023a; Li et al., 2023), while others have focused on the synergy between multilingual NLs and PLs. Notably, ERNIE-Code (Chai et al., 2022) pioneered multilingual LLMs by pre-training across 116 NLs and 6 PLs. Likewise, BLOOM (Scao et al., 2022) was pre-trained on data in 46 NLs and 13 PLs. However, it did not produce competitive results in Python code generation. Recent LLMs like GPT3.5 (OpenAI, 2022) and GPT-4 (OpenAI, 2023) also cater to different NLs and PLs. A significant gap in current code generation benchmarks is the absence of multilingual code generation, preventing an assessment of the capability of LLMs on cross-lingual natural language generalization. Our benchmark aims to address this gap comprehensively.
@@ 4.1.   Experimental Setup
We evaluate the performance of three different LLMs on our benchmark: CodeT5+ (Wang et al., 2023), CodeGen2 (Nijkamp et al., 2023a), two GPT variants, namely GPT-3.5 (OpenAI, 2022) and GPT-4 (OpenAI, 2023). CodeT5+ and CodeGen2 are specialized models designed for code generation tasks, whereas GPT-3.5 and GPT-4 are general-purpose LLMs. For CodeT5+, we experiment with different parameter scales including 220M, 770M and 1B. Similarly, we also evaluate CodeGen2 with 1B, 3.7B, 7B and 16B [^1] Due to page limits, we put all experimental results in appendix and only report different models with largest parameters in the main context, as the observed trends remain consistent across different scales for each model.. Due to constrained computing resources, we report pass@1 for all experiments.
Figure 2: Performance of LLMs, including CodeT5+ (2B), CodeGen2 (3.7B, 16B), GPT-3.5, and GPT-4, on our proposed benchmark. We report pass@1 metric for evaluation. We order languages in their resource availability as summarized in CC100 XL corpus (Lin et al., 2022). The red line in subfigure (h) indicates 0% pass@1.
@@ 4.2.   Results
From Figure 2, we can observe clear disparities in performance across different NLs. This variation underscores the ongoing challenge of achieving cross-lingual NL generalization for current LLMs. Notably, GPT-4 consistently demonstrates superior performance across different PLs and NLs, outperforming all other models significantly. Interestingly, despite the number of parameters of GPT-3.5 is considerably higher than CodeGen2-16B, a specifically tailored code generation modelLLM, GPT-3.5 lags behind CodeGen2-16B in all PLs except Python. This underscores the crucial role of training on coding data.
@@
Python stands out as the most effectively tackled PL by evaluated LLMs. Conversely, all models struggle with JavaScript, C#, Perl, Swift, and TypeScript, as depicted in sub-figure (h). This highlights the formidable complexity of code generation tasks in a multilingual PL context.
@@ Acknowledgments
  • Chai et al. (2022) Yekun Chai, Shuohuan Wang, Chao Pang, Yu Sun, Hao Tian, and Hua Wu. 2022. Ernie-code: Beyond english-centric cross-lingual pretraining for programming languages. In Annual Meeting of the Association for Computational Linguistics.
@@ Appendix B Comprehensive Experiment Results
Chinese 0.00 0.00 0.00 8.75 15.00 21.25 20.00 65.00 78.75
German 0.00 0.00 0.00 8.75 16.25 18.75 21.25 65.00 80.00
Spanish 0.00 1.25 0.00 8.75 15.00 18.75 20.00 63.75 78.75
French 0.00 0.00 0.00 8.75 13.75 21.25 20.00 62.50 77.50
Italian 0.00 0.00 0.00 12.50 15.00 17.50 20.00 70.00 81.25
@@
Bulgarian 0.00 0.00 0.00 8.75 13.75 18.75 18.75 61.25 75.00
Persian 0.00 0.00 0.00 11.25 12.50 16.25 17.50 60.00 78.75
Malay 0.00 0.00 0.00 8.75 13.75 18.75 21.25 62.50 82.50
Hebrew 0.00 0.00 1.25 10.00 13.75 18.75 18.75 60.00 76.25
Estonian 0.00 0.00 0.00 11.25 16.25 16.25 20.00 58.75 81.25
@@
German 0.00 0.00 0.00 8.75 12.50 16.25 20.00 0.00 62.50
Spanish 0.00 0.00 0.00 11.25 11.25 17.50 23.75 0.00 62.50
French 0.00 0.00 0.00 7.50 13.75 18.75 20.00 0.00 57.50
@@
Greek 0.00 1.25 0.00 7.50 11.25 17.50 15.00 0.00 58.75
Hungarian 0.00 0.00 0.00 7.50 16.25 16.25 17.50 0.00 56.25
Dutch 0.00 0.00 0.00 10.00 15.00 16.25 17.50 0.00 63.75
@@
Vietnamese 0.00 0.00 0.00 7.50 11.25 16.25 20.00 0.00 60.00
Bulgarian 0.00 0.00 0.00 10.00 10.00 20.00 18.75 0.00 60.00
Persian 0.00 0.00 0.00 7.50 8.75 16.25 15.00 0.00 58.75
Malay 0.00 0.00 0.00 7.50 11.25 15.00 18.75 0.00 58.75
@@
Estonian 0.00 0.00 0.00 10.00 11.25 13.75 15.00 0.00 55.00
Tagalog 0.00 0.00 0.00 8.75 8.75 12.50 17.50 0.00 70.00
Afrikaans 0.00 0.00 0.00 7.50 12.50 15.00 16.25 0.00 58.75
@@
English 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Russian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Chinese 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
German 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Spanish 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
French 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Italian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Portuguese 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Greek 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Hungarian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Dutch 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Finnish 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Indonesian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Turkish 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Arabic 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Vietnamese 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Bulgarian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Persian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Malay 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Hebrew 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Estonian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Tagalog 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Afrikaans 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
@@
English 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Russian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Chinese 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
German 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Spanish 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
French 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Italian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Portuguese 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Greek 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Hungarian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Dutch 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Finnish 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Indonesian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Turkish 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Arabic 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Vietnamese 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Bulgarian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Persian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Malay 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Hebrew 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Estonian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Tagalog 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Afrikaans 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
@@
English 0.00 0.00 0.00 2.50 10.00 12.50 13.75 2.50 43.75
Russian 0.00 0.00 0.00 3.75 10.00 15.00 15.00 5.00 45.00
Chinese 0.00 0.00 0.00 3.75 10.00 17.50 18.75 2.50 37.50
German 0.00 1.25 0.00 2.50 7.50 11.25 16.25 2.50 40.00
Spanish 0.00 0.00 0.00 3.75 7.50 12.50 13.75 3.75 41.25
French 0.00 0.00 0.00 2.50 7.50 16.25 12.50 2.50 36.25
Italian 0.00 0.00 0.00 3.75 7.50 16.25 15.00 3.75 42.50
Portuguese 0.00 1.25 0.00 2.50 11.25 15.00 16.25 3.75 41.25
Greek 0.00 0.00 0.00 2.50 7.50 11.25 15.00 6.25 28.75
Hungarian 0.00 0.00 0.00 3.75 8.75 13.75 13.75 3.75 43.75
Dutch 0.00 1.25 0.00 3.75 7.50 13.75 15.00 2.50 42.50
Finnish 0.00 1.25 0.00 2.50 7.50 12.50 16.25 1.25 35.00
Indonesian 0.00 1.25 0.00 2.50 7.50 13.75 17.50 1.25 32.50
Turkish 0.00 0.00 0.00 3.75 6.25 11.25 16.25 5.00 45.00
Arabic 0.00 0.00 0.00 3.75 6.25 8.75 15.00 3.75 33.75
Vietnamese 0.00 1.25 0.00 3.75 7.50 13.75 17.50 2.50 40.00
Bulgarian 0.00 0.00 0.00 3.75 7.50 15.00 15.00 7.50 41.25
Persian 0.00 0.00 0.00 2.50 7.50 12.50 13.75 5.00 41.25
Malay 0.00 1.25 0.00 3.75 7.50 11.25 17.50 5.00 33.75
Hebrew 0.00 0.00 0.00 2.50 7.50 12.50 16.25 6.25 35.00
Estonian 0.00 1.25 0.00 5.00 6.25 13.75 17.50 6.25 38.75
Tagalog 0.00 1.25 0.00 2.50 6.25 10.00 15.00 2.50 33.75
Afrikaans 0.00 0.00 0.00 3.75 7.50 11.25 18.75 3.75 36.25
@@
English 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Russian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Chinese 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
German 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Spanish 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
French 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Italian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Portuguese 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Greek 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Hungarian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Dutch 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Finnish 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Indonesian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Turkish 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Arabic 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Vietnamese 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Bulgarian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Persian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Malay 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Hebrew 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Estonian 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Tagalog 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Afrikaans 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00

### 24 Mar 2024 v2

@@ 1.   Introduction
Large language models (LLMs) have made significant progress in code generation, with a predominant focus on the evaluation of generating Python code within an English-centric context (Chen et al., 2021; Austin et al., 2021; Hendrycks et al., 2021; Nijkamp et al., 2023b). In recent years, there has been a growing interest in expanding beyond English-Python code generation to English-multilingual programming languages (PLs) (Athiwaratkun et al., 2023). Some works have also made efforts to incorporate different natural languages (NLs), such as Spanish and Russian, although their coverage of different natural languages remains rather limited (Wang et al., 2022; Chai et al., 2023; Scao et al., 2022). A noticeable gap in these benchmarks is their limited ability to provide a rigorous assessment of multilingual NL to code generation. This overlooks the wide range of different NLs, raising important questions to the evaluation of cross-lingual NL generalization of LLMs. It further underscores the need for a more comprehensive evaluation framework that addresses the gap and captures the diverse linguistic landscape in code generation.
@@ 2.   Related work
LLMs undergo a critical phase of pre-training on a rich collection of resources that include both multilingual PL data and NL data. Some studies have delved into the intricacies of LLMs in multiple PLs (Fried et al., 2023; Wang et al., 2023; Nijkamp et al., 2023a; Li et al., 2023), while others have focused on the synergy between multilingual NLs and PLs. Notably, ERNIE-Code (Chai et al., 2023) pioneered multilingual LLMs by pre-training across 116 NLs and 6 PLs. Likewise, BLOOM (Scao et al., 2022) was pre-trained on data in 46 NLs and 13 PLs. However, it did not produce competitive results in Python code generation. Recent LLMs like GPT3.5 (OpenAI, 2022) and GPT-4 (OpenAI, 2023) also cater to different NLs and PLs. A significant gap in current code generation benchmarks is the absence of multilingual code generation, preventing an assessment of the capability of LLMs on cross-lingual natural language generalization. Our benchmark aims to address this gap comprehensively.
@@ 4.1.   Experimental Setup
We evaluate the performance of three different LLMs on our benchmark: CodeT5+ (Wang et al., 2023), CodeGen2 (Nijkamp et al., 2023a), two GPT variants, namely GPT-3.5 (OpenAI, 2022) and GPT-4 (OpenAI, 2023). CodeT5+ and CodeGen2 are specialized models designed for code generation tasks, whereas GPT-3.5 and GPT-4 are general-purpose LLMs. For CodeT5+, we experiment with different parameter scales including 220M, 770M and 2B. Similarly, we also evaluate CodeGen2 with 1B, 3.7B, 7B and 16B [^1] Due to page limits, we put all experimental results in appendix and only report different models with largest parameters in the main context, as the observed trends remain consistent across different scales for each model.. Due to constrained computing resources, we report pass@1 for all experiments.
Figure 2: Performance of LLMs, including CodeT5+ (2B), CodeGen2 (3.7B, 16B), GPT-3.5, and GPT-4, on our proposed benchmark. We report pass@1 metric for evaluation. We order languages in their resource availability as summarized in CC100 XL corpus (Lin et al., 2022).
@@ 4.2.   Results
From Figure 2, we can observe clear disparities in performance across different NLs. This variation underscores the ongoing challenge of achieving cross-lingual NL generalization for current LLMs. Notably, GPT-4 consistently demonstrates superior performance across different PLs and NLs (except C#), outperforming all other models significantly. Interestingly, despite the number of parameters of GPT-3.5 is considerably higher than CodeGen2-16B, a specifically tailored code generation modelLLM, GPT-3.5 lags behind CodeGen2-16B in all PLs except Python. This underscores the crucial role of training on coding data.
@@
Python stands out as the most effectively tackled PL by evaluated LLMs. Conversely, all models seem to struggle with Scala and Go, as depicted in the figure. This highlights the formidable complexity of code generation tasks in a multilingual PL context.
@@ Acknowledgments
  • Chai et al. (2023) Yekun Chai, Shuohuan Wang, Chao Pang, Yu Sun, Hao Tian, and Hua Wu. 2023. ERNIE-code: Beyond English-centric cross-lingual pretraining for programming languages. In Findings of the Association for Computational Linguistics: ACL 2023, pages 10628–10650, Toronto, Canada. Association for Computational Linguistics.
@@ Appendix B Comprehensive Experiment Results
Chinese 0.00 0.00 1.25 8.75 15.00 21.25 20.00 65.00 78.75
German 0.00 0.00 1.25 8.75 16.25 18.75 21.25 65.00 80.00
Spanish 0.00 1.25 1.25 8.75 15.00 18.75 20.00 63.75 78.75
French 0.00 1.25 1.25 8.75 13.75 21.25 20.00 62.50 77.50
Italian 0.00 0.00 1.25 12.50 15.00 17.50 20.00 70.00 81.25
@@
Bulgarian 0.00 0.00 1.25 8.75 13.75 18.75 18.75 61.25 75.00
Persian 0.00 0.00 1.25 11.25 12.50 16.25 17.50 60.00 78.75
Malay 0.00 0.00 1.25 8.75 13.75 18.75 21.25 62.50 82.50
Hebrew 0.00 1.25 1.25 10.00 13.75 18.75 18.75 60.00 76.25
Estonian 0.00 0.00 1.25 11.25 16.25 16.25 20.00 58.75 81.25
@@
German 0.00 1.25 0.00 8.75 12.50 16.25 20.00 0.00 62.50
Spanish 0.00 1.25 0.00 11.25 11.25 17.50 23.75 0.00 62.50
French 0.00 1.25 0.00 7.50 13.75 18.75 20.00 0.00 57.50
@@
Greek 0.00 2.50 0.00 7.50 11.25 17.50 15.00 0.00 58.75
Hungarian 0.00 1.25 0.00 7.50 16.25 16.25 17.50 0.00 56.25
Dutch 0.00 1.25 0.00 10.00 15.00 16.25 17.50 2.50 63.75
@@
Vietnamese 0.00 1.25 0.00 7.50 11.25 16.25 20.00 1.25 60.00
Bulgarian 0.00 1.25 0.00 10.00 10.00 20.00 18.75 0.00 60.00
Persian 0.00 1.25 0.00 7.50 8.75 16.25 15.00 0.00 61.25
Malay 0.00 1.25 0.00 7.50 11.25 15.00 18.75 1.25 61.25
@@
Estonian 0.00 1.25 0.00 10.00 11.25 13.75 15.00 0.00 55.00
Tagalog 0.00 1.25 0.00 8.75 8.75 12.50 17.50 0.00 70.00
Afrikaans 0.00 1.25 0.00 7.50 12.50 15.00 16.25 0.00 58.75
@@
English 0.00 0.00 0.00 11.25 12.50 17.50 20.00 3.75 80.00
Russian 0.00 2.50 0.00 6.25 17.50 16.25 21.25 2.50 81.25
Chinese 0.00 2.50 0.00 10.00 17.50 17.50 23.75 3.75 76.25
German 0.00 0.00 0.00 6.25 12.50 20.00 20.00 7.50 76.25
Spanish 0.00 2.50 0.00 13.75 12.50 16.25 23.75 8.75 75.00
French 0.00 1.25 0.00 10.00 12.50 21.25 21.25 6.25 72.50
Italian 0.00 1.25 0.00 8.75 12.50 16.25 23.75 8.75 78.75
Portuguese 0.00 2.50 0.00 12.50 12.50 17.50 21.25 8.75 81.25
Greek 0.00 2.50 0.00 7.50 15.00 17.50 20.00 1.25 70.00
Hungarian 0.00 0.00 0.00 10.00 11.25 16.25 21.25 7.50 81.25
Dutch 0.00 0.00 0.00 11.25 17.50 21.25 21.25 8.75 75.00
Finnish 0.00 1.25 0.00 7.50 12.50 15.00 20.00 5.00 77.50
Indonesian 0.00 0.00 0.00 8.75 15.00 17.50 21.25 5.00 75.00
Turkish 0.00 0.00 0.00 8.75 13.75 18.75 21.25 2.50 73.75
Arabic 0.00 2.50 0.00 7.50 15.00 12.50 16.25 1.25 68.75
Vietnamese 0.00 3.75 0.00 12.50 18.75 15.00 22.50 5.00 76.25
Bulgarian 0.00 3.75 0.00 10.00 11.25 17.50 23.75 1.25 78.75
Persian 0.00 3.75 0.00 7.50 16.25 16.25 21.25 3.75 73.75
Malay 0.00 0.00 0.00 8.75 12.50 16.25 25.00 1.25 73.75
Hebrew 0.00 1.25 0.00 8.75 15.00 16.25 18.75 2.50 72.50
Estonian 0.00 0.00 0.00 11.25 11.25 16.25 22.50 2.50 78.75
Tagalog 0.00 0.00 0.00 8.75 15.00 16.25 18.75 6.25 77.50
Afrikaans 0.00 2.50 0.00 8.75 11.25 17.50 21.25 6.25 77.50
@@
English 0.00 3.75 0.00 8.75 10.00 18.75 22.50 2.50 51.25
Russian 0.00 1.25 0.00 6.25 8.75 17.50 18.75 3.75 60.00
Chinese 0.00 1.25 0.00 7.50 12.50 13.75 21.25 3.75 57.50
German 0.00 2.50 0.00 8.75 10.00 13.75 20.00 5.00 62.50
Spanish 0.00 3.75 0.00 10.00 10.00 17.50 22.50 3.75 63.75
French 0.00 2.50 0.00 10.00 11.25 16.25 21.25 3.75 57.50
Italian 0.00 2.50 0.00 8.75 8.75 15.00 25.00 3.75 56.25
Portuguese 0.00 2.50 0.00 10.00 11.25 15.00 23.75 6.25 63.75
Greek 0.00 1.25 0.00 6.25 10.00 12.50 13.75 2.50 53.75
Hungarian 0.00 3.75 0.00 7.50 10.00 12.50 12.50 6.25 57.50
Dutch 0.00 2.50 0.00 7.50 10.00 16.25 22.50 7.50 56.25
Finnish 0.00 1.25 0.00 6.25 10.00 13.75 15.00 2.50 56.25
Indonesian 0.00 2.50 0.00 7.50 10.00 11.25 21.25 2.50 56.25
Turkish 0.00 1.25 0.00 7.50 11.25 13.75 23.75 3.75 55.00
Arabic 0.00 0.00 0.00 6.25 11.25 13.75 21.25 5.00 56.25
Vietnamese 0.00 3.75 0.00 7.50 8.75 13.75 21.25 3.75 58.75
Bulgarian 0.00 5.00 0.00 8.75 11.25 15.00 18.75 3.75 58.75
Persian 0.00 1.25 0.00 6.25 11.25 13.75 16.25 3.75 57.50
Malay 0.00 2.50 0.00 7.50 11.25 12.50 20.00 1.25 60.00
Hebrew 0.00 1.25 0.00 5.00 11.25 13.75 16.25 3.75 53.75
Estonian 0.00 2.50 0.00 6.25 11.25 13.75 17.50 3.75 56.25
Tagalog 0.00 1.25 0.00 7.50 10.00 12.50 17.50 2.50 63.75
Afrikaans 0.00 2.50 0.00 7.50 8.75 13.75 17.50 6.25 52.50
@@
English 0.00 1.25 0.00 2.50 10.00 12.50 13.75 2.50 47.50
Russian 0.00 0.00 0.00 3.75 10.00 15.00 15.00 5.00 50.00
Chinese 0.00 1.25 0.00 3.75 10.00 17.50 18.75 2.50 47.50
German 0.00 1.25 0.00 2.50 7.50 11.25 16.25 2.50 41.25
Spanish 0.00 0.00 0.00 3.75 7.50 12.50 13.75 3.75 45.00
French 0.00 0.00 0.00 2.50 7.50 16.25 12.50 2.50 43.75
Italian 0.00 0.00 0.00 3.75 7.50 16.25 15.00 3.75 48.75
Portuguese 0.00 1.25 0.00 2.50 11.25 15.00 16.25 3.75 45.00
Greek 0.00 0.00 0.00 2.50 7.50 11.25 15.00 6.25 41.25
Hungarian 0.00 0.00 0.00 3.75 8.75 13.75 13.75 3.75 45.00
Dutch 0.00 1.25 0.00 3.75 7.50 13.75 15.00 2.50 50.00
Finnish 0.00 1.25 0.00 2.50 7.50 12.50 16.25 1.25 43.75
Indonesian 0.00 1.25 0.00 2.50 7.50 13.75 17.50 1.25 38.75
Turkish 0.00 1.25 0.00 3.75 6.25 11.25 16.25 5.00 45.00
Arabic 0.00 1.25 0.00 3.75 6.25 8.75 15.00 3.75 43.75
Vietnamese 0.00 1.25 0.00 3.75 7.50 13.75 17.50 2.50 43.75
Bulgarian 0.00 0.00 0.00 3.75 7.50 15.00 15.00 7.50 47.50
Persian 0.00 1.25 0.00 2.50 7.50 12.50 13.75 5.00 45.00
Malay 0.00 1.25 0.00 3.75 7.50 11.25 17.50 5.00 47.50
Hebrew 0.00 0.00 0.00 2.50 7.50 12.50 16.25 6.25 45.00
Estonian 0.00 1.25 0.00 5.00 6.25 13.75 17.50 6.25 43.75
Tagalog 0.00 1.25 0.00 2.50 6.25 10.00 15.00 2.50 40.00
Afrikaans 0.00 0.00 0.00 3.75 7.50 11.25 18.75 3.75 41.25
@@
English 0.00 0.00 0.00 0.00 2.50 1.25 2.50 16.25 62.50
Russian 0.00 0.00 0.00 1.25 0.00 1.25 5.00 16.25 46.25
Chinese 0.00 0.00 0.00 2.50 1.25 1.25 3.75 18.75 52.50
German 0.00 0.00 0.00 1.25 2.50 1.25 1.25 25.00 55.00
Spanish 0.00 0.00 0.00 1.25 1.25 3.75 2.50 25.00 52.50
French 0.00 0.00 0.00 0.00 2.50 2.50 1.25 22.50 48.75
Italian 0.00 0.00 0.00 0.00 0.00 2.50 2.50 26.25 52.50
Portuguese 0.00 0.00 0.00 0.00 2.50 0.00 3.75 21.25 56.25
Greek 0.00 0.00 0.00 1.25 1.25 2.50 0.00 10.00 48.75
Hungarian 0.00 0.00 0.00 1.25 1.25 1.25 0.00 27.50 57.50
Dutch 0.00 0.00 0.00 2.50 1.25 1.25 3.75 26.25 50.00
Finnish 0.00 0.00 0.00 0.00 1.25 1.25 1.25 18.75 52.50
Indonesian 0.00 0.00 0.00 2.50 1.25 0.00 1.25 17.50 55.00
Turkish 0.00 0.00 0.00 0.00 0.00 1.25 2.50 23.75 51.25
Arabic 0.00 0.00 0.00 1.25 1.25 0.00 1.25 18.75 50.00
Vietnamese 0.00 0.00 0.00 1.25 2.50 1.25 1.25 12.50 52.50
Bulgarian 0.00 0.00 0.00 1.25 1.25 2.50 1.25 20.00 50.00
Persian 0.00 0.00 0.00 1.25 0.00 1.25 1.25 23.75 52.50
Malay 0.00 0.00 0.00 1.25 2.50 1.25 1.25 16.25 55.00
Hebrew 0.00 0.00 0.00 0.00 0.00 2.50 3.75 11.25 48.75
Estonian 0.00 0.00 0.00 0.00 0.00 1.25 2.50 21.25 52.50
Tagalog 0.00 0.00 0.00 0.00 0.00 2.50 2.50 18.75 52.50
Afrikaans 0.00 0.00 0.00 1.25 3.75 1.25 1.25 20.00 50.00